At t = 0, a flywheel has an angular velocity of 3.9 rad/s, an angular acceleration of -0.44 rad/s2, and a reference line at θ0 = 0. (a) Through what maximum angle θmax will the reference line turn in the positive direction? What are the (b) first and (c) second times the reference line will be at θ = θmax/4? At what (d) negative time and (e) positive time will the reference line be at θ = -9.9 rad?
Answer #1

(a) Through what maximum angle \thetamax will the reference line turn in the positive direction?

using a kinematic equation 3, we have

\omega2 = \omegai2 + 2 \alpha\thetamax

where, \omega = final velocity = 0 rad/s

THEN, we get

(3.9 rad/s)2 = 2 (0.44 rad/s2) \thetamax

\thetamax = (15.21 rad2/s2) / (0.88 rad/s2)

\thetamax = 17.2 rad

What are the (b) first and (c) second times the reference line will be at \theta = \thetamax /4?

using a kinematic equation 2, we have

\theta = \omegai t + (1/2) \alpha t2

(\thetamax /4) = \omegai t + (0.5) \alpha t2

[(17.2 rad) / 4] = (3.9 rad/s) t - (0.5) (0.44 rad/s2) t2

(0.22 rad/s2) t2 - (3.9 rad/s) t + (4.3 rad) = 0

it's an quadratic equation and we get -

t = 1.18 sec    or     t = 16.5 sec

At what (d) negative time and (e) positive time will the reference line be at \theta = -9.9 rad?

using a kinematic equation 2, we have

\theta = \omegai t + (1/2) \alpha t2

(-9.9 rad) = (3.9 rad/s) t - (0.5) (0.44 rad/s2) t2

(0.22 rad/s2) t2 - (3.9 rad/s) t - (9.9 rad) = 0

it's an quadratic equation and we get -

t = - 2.25 sec    or     t = 19.9 sec